# onlineReferences

### v5

@@ E.4   Revisiting the issue regarding the use of the Wald-DID estimand
$=$ $(\delta_{0,1,AT^{T}}-\delta_{0,0,AT^{T}})\cdot w_{AT^{T},0}+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ in an unexposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (33).
Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows.
@@
$=$ $\Big{\{}\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$-\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}-\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}\Big{\}}$
$+$ $\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1),$
where the type SW (switcher) includes the units who switch from non-treatment to treatment during the two periods. The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (34). Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 27 and Assumption 28, the Wald-DID estimand captures the SLATET under Assumption 29 (parallel trends assumption in the untreated outcome) and conditions (33) and (34). From the above discussion, one can see that imposing the parallel trends assumption in the untreated outcome leads dCDH to assume conditions (33) and (34).
Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET. This also stems from their reliance on the parallel trends assumption in the untreated outcome.
Remark E.4.1.
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 3-4). This identification result, however, is not intuitive because the identification variation comes from a policy shock (instrument) in period $1$.
Remark E.4.2.
dCDH exclude the time compliers $CM^{T}$ in an unexposed group in order to avoid the SLATET in an exposed group being offset by the SLATET’ in an unexposed group. If we add and subtract the average time trends of the untreated outcome for the time compliers, we additionally capture the SLATET’ in an unexposed group from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$.
Thus far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome. Finally, we argue that this assumption would be difficult to rely on in DID-IV settings for three reasons.
[…]
@@
Finally, this assumption is generally not testable using the pre-exposed period data, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
@@
Although dCDH propose the placebo test in their appendix, it has shortcomings for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test has two steps. [^31] Although dCDH ”assume” the conditions (35) and (36), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 29. We, therefore, describe the dCDH’s test in two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
@@
dCDH’s test has some weaknesses. First, the conditions (35) and (36), which should be checked in the first step, are strong, but no guidance is provided when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 29 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$. Indeed, dCDH describe the equality (37) as a necessary condition for Assumption 29 and Assumption 30 to be satisfied under the conditions (35) and (36). Finally, it is not clear how this test should be conducted with multiple pre-exposed period data. Although dCDH assume “for instance that data is available for the period $T=-1$”, it is not indicated how this test can be performed in multiple pre-exposed period settings.
Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (35)-(36) in the first step with the weaker one.
@@
Under repeated cross section settings, the conditions (37)-(39) coincide with the conditions (1)-(2). This indicates that the rationale behind dCDH’s test is more suitable for the assessment of parallel trends assumptions in DID-IV designs.
E.5   Other connections
In this subsection, we document the relationship between DID-IV and Fuzzy DID under non-binary, ordered treatment, and multiple groups settings. We also compare the Wald-DID estimand to the time-corrected Wald estimand considered in dCDH.
E.5.1 Non-binary, ordered treatment
In this subsection, we investigate the relationship between DID-IV to Fuzzy DID under non-binary, ordered treatment settings.
First, we compare the target parameter between the two designs. With a non-binary, ordered treatment $D\in\{0,1,\dots,J\}$, dCDH consider the following target parameter, which we call the switcher average causal response (SACR).
$SACR\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(0)<j<D_{1}(1)),$
@@ E.5.1 Non-binary, ordered treatment
[…]
Next, we compare the identifying assumptions between the two designs. In order for the Wald-DID estimand to identify the SACR, dCDH replace Assumption 28 with the following one.
@@
In addition, dCDH replace the condition (27) in Assumption 26 with the condition (41), and Assumption 27 with the condition (42), respectively.
@@
$D_{01}\sim D_{00},$ (42)
where $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
Note that the condition (42) is stronger than Assumption 27 because this assumption requires that the distribution of the treatment should be stable during the two periods in an unexposed group.
In our DID-IV designs, the identifying assumptions required for the Wald-DID estimand to capture the ACRT are unchanged under non-binary, ordered treatment settings. In addition, Assumptions 17-18 are indirectly testable using the pre-exposed period data in practice.
E.5.2 Multiple groups
In multiple groups settings, our DID-IV design is not directly compatible to Fuzzy DID design because dCDH exploit the group structure in the data ($G\in\{0,1,\dots,\bar{g}\}$) for the Wald-DID estimand to identify the following target parameter.
@@ E.5.2 Multiple groups
where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ represents the switchers in group $G=g$ [^32] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
[…]
In some applications, a researcher might prefer our approach to dCDH’s approach for three reasons. First, our approach yields a policy-relevant parameter by construction if the instrument represents a policy change of interest to the researcher. Second, in multiple period settings, our approach avoids the need to estimate the super groups in each consecutive period $(t-1,t)$, which simplifies both estimation and inference, making our method easier to implement in practice.
Finally, in our approach, the monotonicity assumption can be easy to justify if the instrument is based on a policy shock. In dCDH’s approach, this assumption might be harder to defend in practice because the group variable $G$ typically represents the geographic feature in the data such as municipality, state, or country. [^34] Note that, in multiple groups settings, dCDH also maintain the monotonicity assumption on each group pair $(g,g^{\prime})$ through Assumption 28.
Remark E.4.3 (Multiple groups settings with an ordered treatment).
[…]
E.5.3 Alternative estimand
In this subsection, we compare the Wald-DID estimand to the alternative estimand proposed by dCDH that does not rely on Assumption 29 and 30.353535In this subsection, we do not consider Wald-CIC estimand proposed by dCDH, as we focus on the mean treatment effect in this article.
As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand defined below.
$W_{TC}\equiv\frac{E[Y|Z=1,T=1]-E[Y+(1-D)\delta_{0}+D\delta_{1}|Z=1,T=0]}{E[D|Z=1,T=1]-E[D|Z=1,T=0]},$
where $\delta_{d}=E[Y|Z=0,T=1,D_{1}(0)=d]-E[Y|Z=0,T=0,D_{0}(0)=d]$.
In order for the TC-Wald estimand to identify the SLATET, dCDH consider the following identifying assumption instead of imposing Assumption 29 and 30.
Assumption 35 (Conditional common trends).
For all $d\in\mathcal{S}(D),$
$E[Y(d)|Z=1,T=1,D_{0}(0)=d]-E[Y(d)|Z=1,T=1,D_{0}(0)=d]$
$=$ $E[Y(d)|Z=0,T=1,D_{0}(0)=d]-E[Y(d)|Z=0,T=1,D_{0}(0)=d].$
Note that with a binary treatment, this assumption exploits the observed outcome path in an unexposed groups similar to Assumption 18 in our DID-IV designs (see Table 4).
While the target parameter differs between the two designs, we see that using the Wald-DID estimand under DID-IV designs would be more appealing than using the TC-Wald estimand under Fuzzy DID designs for two reasons. First, when using the Wald-DID estimand in DID-IV designs, researchers can easily assess the plausibility of parallel trends assumptions in practice, employing the pre-trends tests described in Section 5.3. Although the TC-Wald estimand avoids the stable treatment effect assumption under Fuzzy DID designs, it presents challenges in testing the plausibility of Assumption 35 in pre-exposed period data: it is conditional on $D_{0}(0)=d$ and imposes restrictions on the mean evolution of $Y(d)$ for all $d\in\mathcal{S}(D)$.
[…]

### v6

@@ D.4.1 Why the Wald-DID estimand requires Assumption 35 under Fuzzy DID
$=$ $(E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=0,T=0,AT^{T}])$
$\times$ $Pr(AT^{T}|Z=0)+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
@@
$=$ $(E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=1,T=0,AT^{T}])$
$\times$ $Pr(AT^{T}|Z=1)+\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1).$
These decompositions clarify why Assumption 35 becomes necessary. Imposing parallel trends on untreated outcomes requires constructing the average time trends of the untreated potential outcome in both groups (the second term in each conditional expectation). Since these trends are not observed for time always-takers $AT^{T}$, this construction necessarily introduces the first term in the above decompositions, which captures changes in the average treatment effect among $AT^{T}$. For the Wald-DID estimand to identify the SLATET, one must therefore impose stable treatment effects for $AT^{T}$, as in Assumption 35.272727Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term E[Y(1)Y(0)|SW,Z=1,T=1]Pr(SW|Z=1)E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1), which is necessary for identifying the SLATET.
Importantly, this restriction is not intrinsic to the Wald-DID estimand itself, but is a consequence of imposing parallel trends on untreated potential outcomes.
D.4.2 Why the issue does not arise under DID-IV
In contrast to Fuzzy DID, the issue described above does not arise under DID-IV. The key difference is that, in DID-IV designs, the parallel trends assumption is imposed on unexposed outcomes rather than untreated potential outcomes. As a result, the Wald-DID estimand does not require imputing the counterfactual untreated potential outcomes for time always-takers, and hence does not require Assumption 35.
D.5  Which parallel trends assumption is suitable for the DID-IV setting
In Subsection D.4, we showed that differences in the parallel trends (PT) assumption lead to different stances toward the use of the Wald-DID estimand across Fuzzy DID and DID-IV designs. In this subsection, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings in terms of (i) its alignment with the source of identifying variation, (ii) empirical relevance, and (iii) testability.
Which PT assumption aligns with the source of identifying variation
The PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes. This is because the latter relies on variation in treatment, while the former exploits variation in instrument.
Note that in the canonical DID designs, we rely on the PT assumption in untreated outcomes because the units are partitioned into treatment and control groups based on the exposure to the treatment in the second period. In the DID-IV settings, however, the units are partitioned into exposed and unexposed groups based on the exposure to the instrument in the second period. Therefore, imposing the PT assumption on unexposed outcomes would be more natural than imposing the PT assumption on untreated potential outcomes in DID-IV settings.
Empirical relevance
In DID-IV settings, the parallel trends (PT) assumption on untreated potential outcomes is often implausible in practice, while the PT assumption on unexposed outcomes is more appealing. More precisely, because imposing parallel trends on untreated potential outcomes is often not attractive, empirical researchers exploit variation in exposure to an instrument and rely on parallel trends in unexposed outcomes in practice.
A canonical example is Duflo (2001), who study the returns to schooling in Indonesia. The outcome is log annual earnings and the treatment is educational attainment. In this context, imposing PT on untreated potential outcomes would require that the average time trends of log earnings be the same across individuals in the absence of schooling. Given the strong selection into education, such an assumption is difficult to justify empirically. To address this concern, Duflo (2001) exploit the staggered rollout of a school construction program as an instrument for schooling and instead rely on the identifying assumption that, in the absence of the program, the evolution of wages and education across cohorts would not have differed systematically across regions. This identifying restriction corresponds precisely to the parallel trends assumptions on treatment and on unexposed outcomes in DID-IV designs.
Testability
@@ D.5  Which parallel trends assumption is suitable for the DID-IV setting
In DID-IV settings, the PT assumption on unexposed outcomes is indirectly testable using pre-exposed period data as we discussed in Subsection 5.3. By contrast, the PT assumption on untreated potential outcomes is not testable in general. This is because some units may already adopt the treatment before period 0, which prevents a comparison of pre-trends in untreated outcomes between exposed and unexposed groups. From this perspective, the PT assumption on unexposed outcomes is more attractive to empirical researchers.
@@
Although dCDH propose a placebo test in their appendix, its scope for assessing the plausibility of the parallel trends assumption in untreated potential outcomes is somewhat limited. Their procedure consists of two steps. [^28] Although dCDH assume conditions (28) and (29), they note that when either condition fails, the second step can no longer be used to assess Assumption 34. We therefore describe their procedure in two steps. First, we examine whether the share of treated units remain unchanged in both groups from T=1T=-1 to $T=0$:
@@
This testing strategy is informative under the maintained conditions (28) and (29), but its applicability may be limited in practice. First, these first-step conditions are strong, and the procedure does not provide guidance when they are violated. Second, the second step does not directly assess the evolution of the mean untreated outcome between exposed and unexposed groups in pre-exposed periods. Indeed, dCDH interpret (30) as a necessary condition for Assumptions 34 and 35 to hold under (28) and (29). Finally, while the discussion focuses on the case in which data are available for a single pre-exposed period, it is not clear how the procedure should be extended to settings with multiple pre-exposed periods.
Notably, dCDH’s procedure corresponds to our pre-trend test in Subsection 5.3 once the first-step conditions (28)–(29) are replaced by the weaker condition:
@@
Under repeated cross-section settings, conditions (30)–(31) coincide with conditions (1)–(2) in Subsection 5.3. This observation suggests that the underlying logic of dCDH’s placebo test is more aligned with the assessment of parallel trends assumptions in DID-IV designs.
D.6  Other comparisons
In this subsection, we document the relationship between DID-IV and Fuzzy DID in settings with non-binary, ordered treatments and multiple groups. We also discuss the alternative estimands proposed in dCDH that avoid Assumption 35.
D.6.1 Non-binary, ordered treatment
Again, we first document the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter and the identifying assumptions.
Empirical setting
In the case of non-binary, ordered treatment, the condition (ii) of the DID-IV setting is satisfied in dCDH because the group variable $G$ is included in their ordered treatment equation (see Assumption 3’ in dCDH). By contrast, to the best of our knowledge, the condition (i) of the DID-IV setting is not imposed in dCDH: the threshold $v^{d}_{GT}$ in their ordered treatment equation is allowed to depend on $G$ in period $T=0$ for all $d\in\{0,\dots,\bar{d}\}$. This implies that the DID-IV setting differs from the setting considered in dCDH.
Nevertheless, based on their empirical application, it is reasonable to infer that dCDH have the DID-IV setting in mind. In particular, dCDH revisit Duflo (2001), in which the outcome $Y$ is log annual earnings, the treatment $D$ is education attainment (an ordered treatment), and the group variable $G\in\{0,1\}$ indicates the regions that are highly exposed to the policy shock in the second period. Given this, the absence of condition (i) in dCDH can be interpreted as reflecting an additional restriction on treatment adoption behavior in period $T=0$.
From the above discussion, in what follows, we replace $G$ with $Z$.
The target parameter
In this paper, we define the ACRT as our target parameter. By contrast, dCDH consider the following target parameter, which we call the switcher average causal response on the treated (SACRT).
$SACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(1)<j<D_{1}(1)),$
@@ D.6.1 Non-binary, ordered treatment
Here, to ease the interpretation, we slightly modify the notation in dCDH. Specifically, we define $D_{t}(Z)=\sum_{j=1}^{J}\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH (see Assumption 39 below).
The estimand
Both papers consider the same Wald-DID estimand. This paper regards the Wald-DID estimand as the natural estimand for identifying the ACRT. By contrast, dCDH point out that this estimand identifies the SACRT only if the stable treatment effect assumption (see Assumption 40 below) is satisfied.
The identifying assumptions
In this paper, we consider the same identifying assumptions under non-binary, ordered treatment settings. By contrast, while dCDH maintain Assumption 31 and Assumption 34, they replace Assumption 33 with the following one.
@@
In addition, they replace Assumption 32 with the following two conditions (33)-(34).
@@
$D_{01}\sim D_{00}.$ (34)
Here, $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution. Under non-binary, ordered treatment settings, Assumption 35 is formulated as follows:
Assumption 40 (Stable treatment effect over time with non-binary, ordered treatment).
For all $d\in\{0,1,\dots,J\}$,
$E[Y(d)-Y(0)|Z,T=1,D_{0}(Z)=d]=E[Y(d)-Y(0)|Z,T=0,D_{0}(Z)=d].$
Next, we compare the restrictions on treatment adoption behavior between this paper and dCDH. We also discuss the difference in target parameter. Finally, we discuss the different stances regarding the use of the Wald-DID estimand.
Restrictions on treatment adoption behavior
In DID-IV, we restrict the treatment adoption behavior through the monotonicity with respect to the instrument in the second period (Assumption 20). [^29] Note that dCDH do not impose this restriction, as Assumption 39 does not imply Assumption 20 in general. However, Assumption 39 implies the following restrictions: $v^{j}_{1t}\geq v^{j}_{0t}\hskip 5.69054pt\text{or}\hskip 5.69054ptv^{j}_{1t}\leq v^{j}_{0t}.$ for all $j\in\{1,\dots,J\}$ and all $t\in\{0,1\}$.
In Fuzzy DID, dCDH impose the restrictions on treatment adoption behavior through the conditions (33)-(34). Specifically, as dCDH point out in their proof (see page 1025 in dCDH), these conditions combined with Assumption 39 imply:
$v^{j}_{01}$ $=v^{j}_{00},$ (35)
$v^{j}_{11}$ $\leq v^{j}_{10}$ (36)
for all $j\in\{1,\dots,J\}$. Here, the condition (35) requires that the treatment rate should be stable over time for all units in the absence of exposure to the instrument. The condition (36) requires that the treatment should be weakly increasing over time for all units under exposure to the instrument.
The difference in target parameter
In general, the SACRT considered in dCDH differs from the ACRT defined in this paper. As in the case of the SLATET, the SACRT may not be policy relevant. This is because the type $SW^{{}^{\prime}}=\{Z=1,T=1,D_{0}(1)<j<D_{1}(1)\}$ is induced to increase the treatment by time, and not equal to the type $\{Z=1,T=1,D_{1}(0)<j<D_{1}(1)\}$ in general.
Note that our target parameter, the ACRT, is the conditional version of the ACR defined in Angrist and Imbens (1995). As a result, the ACRT is policy-relevant parameter by construction.
The use of the Wald-DID estimand
dCDH point out that the stable treatment effect assumption (Assumption 40) is also required for the Wald–DID estimand to identify the SACRT. As in the binary treatment case, however, this requirement arises from imposing the parallel trends assumption on untreated outcomes.
To see this formally, we first rewrite Assumption 40 as follows:
$E[Y(d)\mid Z,T=1,D_{0}(Z)=d]-E[Y(d)\mid Z,T=0,D_{0}(Z)=d]$
$=$ $E[Y(0)\mid Z,T=1,D_{0}(Z)=d]-E[Y(0)\mid Z,T=0,D_{0}(Z)=d].$
From the above expression, this assumption can be interpreted as requiring that, for units with $D_{0}(Z)=d$, the average time trend of the potential outcome $Y(d)$ coincides with that of the untreated potential outcome $Y(0)$, conditional on $Z$.
Recall that under Fuzzy DID with ordered treatments, Assumptions (33)–(34) imply that treatment adoption is stable over time in the unexposed group and weakly increasing over time in the exposed group. As a result, units with $D_{0}(0)=d$ in the unexposed group necessarily remain at treatment level $d$ under $T=1$, while in the exposed group such units can be partitioned into those whose treatment remains at $d$ and those whose treatment increases from $d$ to $d^{\prime}$.
For units whose treatment remains constant over time, the left-hand side of Assumption 40 is observed. The assumption therefore uses this observed quantity to impute the unobserved average time trends of untreated potential outcomes for these units. For units in the exposed group whose treatment increases from $d$ to $d^{\prime}$, the same assumption additionally allows one to isolate the causal effect associated with the change from $d$ to $d^{\prime}$, which constitutes a component of the SACRT.
In short, Assumption 40 is required not because of the Wald–DID estimand itself, but because imposing the parallel trends assumption on untreated outcomes forces one to impute the unobserved evolution of untreated potential outcomes for units with $D_{0}(Z)=d$.
D.6.2 Multiple groups
In multiple group settings, the DID-IV design considered in this paper is not directly compatible with the Fuzzy DID framework. There are two reasons. First, the group variable $G$ takes a finite number of ordered values: $G\in\{0,1,\dots,\bar{g}\}$. Second, dCDH construct the “super groups” (see dCDH) from the group structure in the data, $G\in\{0,1,\dots,\bar{g}\}$, and use the Wald-DID estimand to identify the following target parameter:
@@ D.6.2 Multiple groups
where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ denotes the set of switchers in group G=gG=g. [^30] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
However, even in multiple group settings, dCDH continue to impose Assumption 33. This implies that their framework still maintains two restrictions on treatment adoption behavior: monotonicity with respect to time $T$ and monotonicity with respect to the group variable $G$, in the sense of Imbens and Angrist (1994).
Remark 5.
While dCDH and its supplemental appendix study multiple groups and ordered treatments separately, to the best of our knowledge, we are not aware of the identification results for settings with multiple groups under ordered treatments in either the main text or the supplement.
D.6.3 Alternative estimand
As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand and the changes-in-changes Wald (CIC-Wald) estimand, which do not rely on Assumptions 34 and 35. Under both estimands, however, dCDH continue to impose Assumptions 31-33 to identify the SLATET. This implies that the same discussions in Subsections D.2-D.3 can be applied to both estimands regarding the restrictions on treatment adoption behavior and the target parameter, the SLATET (or SACRT). These estimands do not require Assumption 35 because Assumption 34 is not imposed for the time-always takers $AT^{T}$.
D.7  Summary and guidance
Throughout this section, we compare the DID-IV with Fuzzy DID considered in dCDH. To sum up, we obtain the following implications. First, the restrictions on treatment adoption behavior under Fuzzy DID are generally stronger than those under DID-IV. Second, the target parameter in Fuzzy DID, the SLATET, is not policy relevant in general, and even when it is policy relevant, it is defined on a strictly smaller population than the LATET. Finally, the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH imposes the parallel trends assumption on untreated outcomes.
We conclude this section by summarizing practical considerations that may guide empirical researchers in choosing between DID-IV and Fuzzy DID designs under the DID-IV setting.
Policy-relevant parameter
If empirical researchers are interested in identifying a policy-relevant parameter, DID-IV designs are preferable. Under Fuzzy DID designs, the target parameter may not be policy relevant in general. If researchers instead focus on the SLATET (or the SACRT), it is recommended that they clarify why the causal effect for units whose treatment status is affected by time but not by the instrument is of substantive interest in their empirical context.
Restrictions on treatment adoption behavior
If empirical researchers are interested in identifying the SLATET (SACRT), they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed under Fuzzy DID designs. If these restrictions appear implausible, researchers may instead target the LATET (ACRT) and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument.
Testability of the parallel trends assumption
If empirical researchers wish to assess the plausibility of the parallel trends assumption, DID-IV designs are preferable. Under Fuzzy DID designs, Assumption 34 (and its counterparts for the alternative estimands) is typically difficult to assess using pre-exposure period data (see Subsection 1.1 of the Supplemental Appendix in dCDH).